OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 7) and (4, 9)? slope = -1/3 slope = 1/3 slope = 3

OpenStudy (anonymous):

would u use the formula y2-y1/x2-x1

OpenStudy (anonymous):

how do u put the numbers in there

OpenStudy (phi):

the formula means find the difference between the y values (the 2nd number in the pair) in (-2,7) 7 is the y value. y2 means the value from the 2nd point. y1 means the y value of the 1st point. pick on of the points to be 1st, and the other to be second. (it does not matter which)

OpenStudy (phi):

then find the difference between the x values.

OpenStudy (anonymous):

so whould the answer be -1/3

OpenStudy (phi):

can you show your work ?

OpenStudy (anonymous):

7-(-2)=9 right

OpenStudy (anonymous):

9-4=5

OpenStudy (phi):

when you have time, this goes over how to do it http://www.khanacademy.org/math/trigonometry/graphs/slope_of_line/v/slope-of-a-line-2 meanwhile, let the 1st point be (-2, 7) and the 2nd point be (4, 9) what is y2 and what is y1 ?

OpenStudy (phi):

remember, each point has only 1 y-value, the 2nd number.

OpenStudy (anonymous):

y2=7 y1=4

OpenStudy (phi):

y2 is the y value of the second point. The second point is (4, 9) what is y2 ?

OpenStudy (phi):

y2 is the 2nd number in (4,9).

OpenStudy (anonymous):

9

OpenStudy (phi):

now what is y1. y1 is the 2nd number in the first point.

OpenStudy (anonymous):

7

OpenStudy (phi):

y2 -y1 is ?

OpenStudy (anonymous):

2

OpenStudy (phi):

ok now you have y2-y1 we need x2 - x1 x2 is the x value of the second point. the x value is the first number in the pair what is x2 ?

OpenStudy (phi):

the 2nd point is (4, 9) what is x2 ?

OpenStudy (anonymous):

9-(-2)

OpenStudy (phi):

don't jump ahead. we need the x value of the second point. the second point is (4,9) what is the first number in (4,9) ? that is the x value?

OpenStudy (anonymous):

4

OpenStudy (phi):

yes, x2 is 4 now what is x1 ? to find x1, look at the first point (-2, 7) find the x value of that point

OpenStudy (anonymous):

Can someone help me with my question?

OpenStudy (anonymous):

-2 so it would be 4-2

OpenStudy (phi):

almost, it would be 4 - -2

OpenStudy (anonymous):

-.- Rude Much!

OpenStudy (anonymous):

so it would be 6

OpenStudy (phi):

yes you have y2-y1= 2 and x2-x1= 6 now we find change in y divided by change in x 2/6 can you simplify that ?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

thank u so much